Example for seminar.sty

Policarpa Salabarrieta July 21, 1991

Information overload = “Too much” information

You have 134 unread messages:
Do you want to read them now?

1. People cannot process all the information they receive.

2. People should receive less information.

• There is information overload in a network if there is some mechanism that, compared to the status quo, makes the senders and/or receivers better off by restricting the flow of information.’

• There is information overload in a network if there is some mechanism that, compared to the status quo, makes the senders and/or receivers better off by restricting the flow of information.

Questions

• When could there be overload in networks?

• What mechanims make the receivers and senders better off?

• How does the welfare of the senders and receivers depend on the cost of communication?

Being more informed

is always better, but it’s not the same as

receiving more information

A tax τ on communication is said to support X̃(c) if X̃(c) is an equilibrium for Γ(c + τ).

Proposition 6. Assume X̃(c) is not an equilibrium for Γ(c).

1. If supp(γ) = [0, 1]^{n}, there is no tax that supports X̃(c).

2. If supp(γ) = S^{n−1}, there is a tax that supports X̃(c) if and only if m = 1, p_{j} > c ∀j, and

(a) n = 2; or

(b) n = 3 and p_{i}^{−1} + p_{j}^{−1} ≥ p_{k}^{−1} for all distinct i, j, k; or

(c) n = 4 and p_{1} = p_{2} = p_{3} = p_{4}.

Architecture

[figure: 1; 2; 3; 4; 5; 6; 7; 8]